# Backbone probability of planar Brownian motion

### v1

@@ 3 The exact laws of the conformal radii
By applying Schwartz’s lemma for $g^{-1}$, we have $g(\frac{1}{2}\mathbbm{D}\cap D(\ell))\supset\frac{1}{2}\mathbbm{D}$, and $\frac{1}{|g^{\prime}(0)|}=\mathrm{CR}(\ell,0)\geq\mathrm{dist}(\ell,0)>5\sqrt{\varepsilon}$. Let $\widetilde{g}(z):=\frac{|g^{\prime}(0)|}{5\sqrt{\varepsilon}}g^{-1}(5\sqrt{\varepsilon}z)$; note that $\widetilde{g}(0)=0$ and $|\widetilde{g}^{\prime}(0)|=1$. By Koebe’s 1/4 theorem, we have $\frac{1}{4}\mathbbm{D}\subset\widetilde{g}(\mathbbm{D})$, which implies $\frac{1}{4}\mathbbm{D}\subset\frac{|g^{\prime}(0)|}{5\sqrt{\varepsilon}}g^{-1}(5\sqrt{\varepsilon}\mathbbm{D})\subset\frac{1}{5\varepsilon}g^{-1}(5\sqrt{\varepsilon}\mathbbm{D})$, as desired. ∎
The upper bound of $\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]$ then relies on the following iterative inequality.
Proposition 4.4.
Let $q_{0}=\mathbbm{P}[\widetilde{\ell}_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset]\in(0,1)$. Then there is a constant $C_{0}>0$ such that for any $\varepsilon\in(0,10^{-4})$,
$\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]\leq\frac{C_{0}}{\log|\log\varepsilon|}+q_{0}\mathbbm{P}[\mathsf{Bac}_{5\sqrt{\varepsilon}}].$
Proof.
Note that
$\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]$ $\leq\mathbbm{P}[\ell_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}\neq\emptyset]+\mathbbm{P}[\ell_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}=\emptyset,\mathsf{Bac}_{\varepsilon}]$
$\leq\mathbbm{P}[\ell_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}\neq\emptyset]+\mathbbm{P}[\widetilde{\ell}_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}\neq\emptyset]+\mathbbm{P}[\widetilde{\ell}_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}=\emptyset,\mathsf{Bac}_{\varepsilon}]$
$=\mathbbm{P}[\ell_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}\neq\emptyset]+\mathbbm{P}[\widetilde{\ell}_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}\neq\emptyset]+\mathbbm{P}[\widetilde{\ell}_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}=\emptyset,\widetilde{\ell}_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset,\mathsf{Bac}_{\varepsilon}]$ (4.1)
Here the last line follows from that $(\widetilde{\ell}_{1}\cap(5\sqrt{\varepsilon}\mathbbm{D}\cup\mathbbm{A}_{1/2})=\emptyset)\cap\mathsf{Bac}_{\varepsilon}=\emptyset$. By Lemma 3.3, Corollary 3.7 and using Koebe’s 1/4 theorem, there exist some positive constants $C,C^{\prime}$ and $C_{0}$ such that
$\mathbbm{P}[\ell_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}\neq\emptyset]+\mathbbm{P}[\widetilde{\ell}_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}\neq\emptyset]$ $\leq\mathbbm{P}[\mathrm{CR}(\ell_{1},0)<20\sqrt{\varepsilon}]+\mathbbm{P}[\mathrm{CR}(\widetilde{\ell}_{1},0)<20\sqrt{\varepsilon}]$
$\leq C\varepsilon^{\frac{1}{8}}+C^{\prime}\frac{1}{\log|\log\varepsilon|}\leq\frac{C_{0}}{\log|\log\varepsilon|}.$ (4.2)
Let $q_{0}:=\mathbbm{P}[\widetilde{\ell}_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset]\in(0,1)$. Then
$\mathbbm{P}[\widetilde{\ell}_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}=\emptyset,\widetilde{\ell}_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset,\mathsf{Bac}_{\varepsilon}]\leq q_{0}\mathbbm{P}[\mathsf{Bac}_{\varepsilon}|\widetilde{\ell}_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}=\emptyset,\widetilde{\ell}_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset].$ (4.3)
Let $g:D(\widetilde{\ell}_{1})\to\mathbbm{D}$ be the conformal map such that $g(0)=0,g(B_{s_{1}})=1$. By Lemma 3.5, we see that $(g(B_{t}))_{t\in[0,s_{1}]}$ is independent of $B[s_{1},\tau_{\mathbbm{D}}]$, and has the same law (up to a time change) as $B[0,\tau_{\mathbbm{D}}]$ conditioned on $B_{\tau_{\mathbbm{D}}}=1$. Hence, if we denote $\widetilde{\mathbbm{P}}$ to be the law of $(g(B_{t}))_{t\in[0,s_{1}]}$ conditioned on the event $\{\widetilde{\ell}_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}=\emptyset\}\cap\{\widetilde{\ell}_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset\}$, then $\widetilde{\mathbbm{P}}$ is still the same as (up to a time change) the law of $B[0,\tau_{\mathbbm{D}}]$ conditioned on $B_{\tau_{\mathbbm{D}}}=1$. By Lemma 4.3, on the event $\{\widetilde{\ell}_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}=\emptyset\}\cap\{\widetilde{\ell}_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset\}$, we have $\overline{g(\varepsilon\mathbbm{D})}\subset 5\sqrt{\varepsilon}\mathbbm{D}$ and $g(\frac{1}{2}\mathbbm{D}\cap D(\widetilde{\ell}_{1}))\supset\frac{1}{2}\mathbbm{D}$. Then we find
$\mathbbm{P}[\mathsf{Bac}_{\varepsilon}|\widetilde{\ell}_{1}\cap 5\sqrt{\varepsilon}\mathbbm{D}=\emptyset,\widetilde{\ell}_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset]\leq\widetilde{\mathbbm{P}}[\mathsf{Bac}_{5\sqrt{\varepsilon}}]=\mathbbm{P}[\mathsf{Bac}_{5\sqrt{\varepsilon}}],$ (4.4)
where the last equality follows from the rotational invariance of $B[0,\tau_{\mathbbm{D}}]$. The result then follows from combining (4.1), (4.2), (4.3) and (4.4). ∎
Now we prove the upper bound for $\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]$, finishing the proof of Theorem 1.1.
@@
Let $\varepsilon\in(0,10^{-4})$ and $F(\varepsilon):=\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]$. Note that the function $x\mapsto\frac{1}{\log|\log x|}$ is increasing on $(0,10^{-4})$ and there is a constant $C>0$ such that $\frac{1}{\log|\log(25x)|}\leq\frac{C}{\log|\log x|}$ for any $x\in(0,10^{-4})$. Then for each $n\in\mathbbm{Z}_{+}$ with $25\varepsilon^{\frac{1}{2^{n}}}<10^{-4}$, by applying Proposition 4.4 for $n$ times, we find
$F(\varepsilon)$ $\leq C_{0}\sum\limits_{k=0}^{n-1}\frac{q_{0}^{k}}{\log|\log(5^{2-\frac{1}{2^{k-1}}}\varepsilon^{\frac{1}{2^{k}}})|}+q_{0}^{n}F(5^{2-\frac{1}{2^{n-1}}}\varepsilon^{\frac{1}{2^{n}}})$
$\leq\frac{C_{0}}{1-q_{0}}\frac{1}{\log|\log(25\varepsilon^{\frac{1}{2^{n}}})|}+q_{0}^{n}\leq\frac{C_{0}C}{1-q_{0}}\frac{1}{\log|\log(\varepsilon^{\frac{1}{2^{n}}})|}+q_{0}^{n}$
$=\frac{C_{0}C}{1-q_{0}}\frac{1}{\log|\log(\varepsilon)|-n\log 2}+q_{0}^{n}.$ (4.5)
Now we choose some proper $n$ in (4.5) to obtain the result. Let
$C^{\prime}:=\max\left\{C,\sup\limits_{\varepsilon\in(0,10^{-4})}\left(\log\log|\log\varepsilon|-\frac{|\log q_{0}|}{\log 2}\left(\log|\log\varepsilon|-\log\log(250000)-\log 2\right)\right)\right\},$
and let $n_{\varepsilon}:=\inf\{n\geq 0:q_{0}^{n}<\frac{C^{\prime}}{\log|\log\varepsilon|}\}$. Note that $n_{\varepsilon}\in\left(\frac{\log\log|\log\varepsilon|-\log C^{\prime}}{|\log q_{0}|},1+\frac{\log\log|\log\varepsilon|-\log C^{\prime}}{|\log q_{0}|}\right)$ and $25\varepsilon^{\frac{1}{2^{n_{\varepsilon}}}}<10^{-4}$. By taking $n=n_{\varepsilon}$ in (4.5), we obtain
$F(\varepsilon)$ $\leq\frac{C_{0}C^{\prime}}{1-q_{0}}\frac{1}{\log|\log(\varepsilon)|-n\log 2}+q_{0}^{n}$
$\leq\frac{C_{0}C^{\prime}}{1-q_{0}}\frac{1}{\log|\log(\varepsilon)|-\frac{\log 2}{|\log q_{0}|}(\log\log|\log\varepsilon|-\log C^{\prime})}+\frac{C^{\prime}}{\log|\log\varepsilon|}\leq\frac{C_{2}}{\log|\log\varepsilon|}$
for some positive constant $C_{2}$, as desired. ∎

### v2

@@ 3 Exact laws of the conformal radii in $B[0,\tau_{\mathbbm{D}}]$
By applying Schwartz’s lemma for $f^{-1}$, we have $|f^{\prime}(0)|^{-1}=\mathrm{CR}(\ell,0)\geq\mathrm{dist}(\ell,0)>2\sqrt{\varepsilon}$. Let $\widetilde{f}(z):=\frac{|f^{\prime}(0)|}{2\sqrt{\varepsilon}}f^{-1}(2\sqrt{\varepsilon}z)$; note that $\widetilde{f}(0)=0$ and $|\widetilde{f}^{\prime}(0)|=1$. By Koebe’s 1/4 theorem, we have $\frac{1}{4}\mathbbm{D}\subset\widetilde{f}(\mathbbm{D})$, which implies $\frac{1}{4}\mathbbm{D}\subset\frac{|f^{\prime}(0)|}{2\sqrt{\varepsilon}}f^{-1}(2\sqrt{\varepsilon}\mathbbm{D})\subset\frac{1}{4\varepsilon}f^{-1}(2\sqrt{\varepsilon}\mathbbm{D})$, as desired. ∎
@@
For $\varepsilon\in(0,10^{-10})$, let $\delta=2\sqrt{\varepsilon}$. Let $A_{n,\delta}=\{\ell_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset,\widetilde{\ell}_{n}\cap\delta\mathbbm{D}\neq\emptyset\}$ be as above. Then by the definition of $A_{1,\delta}$, we have
$\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]$ $\leq\mathbbm{P}[\mathsf{Bac}_{\varepsilon}\backslash\cup_{n\geq 1}A_{n,\delta}]+\mathbbm{P}[\cup_{n\geq 1}A_{n,\delta}]$
[…]
Here the last inequality uses the fact that $\{\mathsf{Bac}_{\varepsilon},\ell_{1}\cap\mathbbm{A}_{1/2}=\emptyset\}$ implies $\ell_{1}\cap\varepsilon\mathbbm{D}\neq\emptyset$, while $\mathsf{Bac}_{\varepsilon}\backslash\cup_{n\geq 2}A_{n,\delta},\ell_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset,\widetilde{\ell}_{1}\cap\delta\mathbbm{D}=\emptyset\}$ implies $\mathsf{Bac}_{\varepsilon}\backslash\cup_{n\geq 2}A_{n,\delta},\widetilde{\ell}_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset,\widetilde{\ell}_{1}\cap\delta\mathbbm{D}=\emptyset\}$. Similarly, by (4.7) and using the definition of $A_{n,\delta}$ iteratively, we have
$\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]$ $\leq\mathbbm{P}[\ell_{1}\cap\varepsilon\mathbbm{D}\neq\emptyset]+\mathbbm{P}[\widetilde{\ell}_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset,\widetilde{\ell}_{1}\cap\delta\mathbbm{D}=\emptyset,\ell_{2}\cap\varepsilon\mathbbm{D}\neq\emptyset]$
$+\mathbbm{P}[\mathsf{Bac}_{\varepsilon}\backslash\cup_{n\geq 3}A_{n,\delta},\widetilde{\ell}_{2}\cap\mathbbm{A}_{1/2}\neq\emptyset,\widetilde{\ell}_{2}\cap\delta\mathbbm{D}=\emptyset]+\mathbbm{P}[\cup_{n\geq 1}A_{n,\delta}]$
$\leq$.
$\leq\mathbbm{P}[\ell_{1}\cap\varepsilon\mathbbm{D}\neq\emptyset]+\sum_{n=1}^{m}\mathbbm{P}[\widetilde{\ell}_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset,\widetilde{\ell}_{n}\cap\delta\mathbbm{D}=\emptyset,\ell_{n+1}\cap\varepsilon\mathbbm{D}\neq\emptyset]$
$+\mathbbm{P}[\mathsf{Bac}_{\varepsilon}\backslash\cup_{n\geq m+2}A_{n,\delta},\widetilde{\ell}_{m+1}\cap\mathbbm{A}_{1/2}\neq\emptyset,\widetilde{\ell}_{m+1}\cap\delta\mathbbm{D}=\emptyset]+\mathbbm{P}[\cup_{n\geq 1}A_{n,\delta}]$
for any $m\geq 1$. Taking $m\to\infty$ and using that $\lim\limits_{n\to\infty}\mathbbm{P}[\widetilde{\ell}_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset]=0$, we obtain that
$\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]\leq\mathbbm{P}[\ell_{1}\cap\varepsilon\mathbbm{D}\neq\emptyset]+\sum_{n\geq 1}\mathbbm{P}[\widetilde{\ell}_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset,\widetilde{\ell}_{n}\cap\delta\mathbbm{D}=\emptyset,\ell_{n+1}\cap\varepsilon\mathbbm{D}\neq\emptyset]+\mathbbm{P}[\cup_{n\geq 1}A_{n,\delta}].$ (4.8)
It suffices to bound each of the three terms on the right side of (4.8). First, by Lemma 3.3 and Koebe’s 1/4 theorem, there exists $C_{1}\in(0,\infty)$ such that
$\mathbbm{P}[\ell_{1}\cap\varepsilon\mathbbm{D}\neq\emptyset]\leq C_{1}\varepsilon^{\frac{1}{4}}.$ (4.9)
Recall that $\widetilde{g}_{n}$ is the conformal map from $D(\widetilde{\ell}_{n})$ to $\mathbbm{D}$ that fixes the origin with $\widetilde{g}_{n}(B_{s_{n}})=1$. By Lemma 4.4, $\{\widetilde{\ell}_{n}\cap\delta\mathbbm{D}=\emptyset\}$ implies $\{\widetilde{g}_{n}(\varepsilon\mathbbm{D})\subset\delta\mathbbm{D}\}$. Furthermore, by Lemma 3.5, we know that $\widetilde{g}_{n}(\ell_{n+1})$ is independent of $\widetilde{\ell}_{n}$, and $\mathbbm{P}[\widetilde{g}_{n}(\ell_{n+1})\cap\delta\mathbbm{D}\neq\emptyset]=\mathbbm{P}[\ell_{1}\cap\delta\mathbbm{D}\neq\emptyset]$. Hence, we have
$\sum_{n\geq 1}\mathbbm{P}[\widetilde{\ell}_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset,\widetilde{\ell}_{n}\cap\delta\mathbbm{D}=\emptyset,\ell_{n+1}\cap\varepsilon\mathbbm{D}\neq\emptyset]$ $\leq\sum_{n\geq 1}\mathbbm{P}[\widetilde{\ell}_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset,\widetilde{\ell}_{n}\cap\delta\mathbbm{D}=\emptyset,\widetilde{g}_{n}(\ell_{n+1})\cap\delta\mathbbm{D}\neq\emptyset]$
$\leq\mathbbm{P}[\ell_{1}\cap\delta\mathbbm{D}\neq\emptyset]\cdot\sum_{n\geq 1}\mathbbm{P}[\widetilde{\ell}_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset]$
$\leq 2C_{1}S\varepsilon^{\frac{1}{8}},$ (4.10)
where the last line uses (4.9) and $S$ is defined in (4.1).
Now let $N_{\delta}=\inf\{n\geq 1:A_{n,\delta}\text{ occur}\}$, and hence $\{\cup_{n\geq 1}A_{n,\delta}\}=\{N_{\delta}<\infty\}$. By Koebe’s 1/4 theorem, we have
$\mathbbm{P}[\cup_{n\geq 1}A_{n,\delta}]$ $=\mathbbm{P}[N_{\delta}<\infty]=\mathbbm{P}[N_{\delta}<\infty,\mathrm{CR}(\widetilde{\ell}_{N_{\delta}},0)\leq 4\delta]$
$\leq\mathbbm{P}\left[N_{\delta}<\infty,\frac{\mathrm{CR}(\widetilde{\ell}_{N_{\delta}},0)}{\mathrm{CR}(\ell_{N_{\delta}},0)}<2\sqrt{\delta}\right]+\mathbbm{P}[N_{\delta}<\infty,\mathrm{CR}(\ell_{N_{\delta}},0)<2\sqrt{\delta}].$ (4.11)
Note that the definition of $N_{\delta}$ implies
$\mathbbm{P}\left[N_{\delta}<\infty,\frac{\mathrm{CR}(\widetilde{\ell}_{N_{\delta}},0)}{\mathrm{CR}(\ell_{N_{\delta}},0)}<2\sqrt{\delta}\right]$ $=\sum\limits_{n=1}^{\infty}\mathbbm{P}\left[N_{\delta}=n,\frac{\mathrm{CR}(\widetilde{\ell}_{n},0)}{\mathrm{CR}(\ell_{n},0)}<2\sqrt{\delta}\right]$
$\leq\sum\limits_{n=1}^{\infty}\mathbbm{P}\left[\ell_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset,\frac{\mathrm{CR}(\widetilde{\ell}_{n},0)}{\mathrm{CR}(\ell_{n},0)}<2\sqrt{\delta}\right]$
$=\mathbbm{P}[\mathrm{CR}(\rho,0)<2\sqrt{\delta}]\cdot S,$
where the last line follows from Lemma 3.5. Meanwhile, $\{N_{\delta}<\infty,\mathrm{CR}(\ell_{N_{\delta}},0)<2\sqrt{\delta}\}$ implies $\{N_{\delta}<\infty,\ell_{N_{\delta}}\cap\mathbbm{A}_{1/2}\neq\emptyset,\ell_{N_{\delta}}\cap 2\sqrt{\delta}\mathbbm{D}\neq\emptyset\}$, which further implies the Brownian trajectory $B[\sigma_{2\sqrt{\delta}},\tau_{1/2}]$ does not disconnect $2\sqrt{\delta}\mathbbm{S}^{1}$ and $\frac{1}{2}\mathbbm{S}^{1}$. Combined with (4.11) and using again the Brownian disconnection exponent $\frac{1}{4}$ from [LSW01b], we find
$\mathbbm{P}[\cup_{n\geq 1}A_{n,\delta}]$ $\leq\mathbbm{P}[\mathrm{CR}(\rho,0)<2\sqrt{\delta}]\cdot S+\mathbbm{P}\left[B[\sigma_{2\sqrt{\delta}},\tau_{1/2}]\text{ does not disconnect }2\sqrt{\delta}\mathbbm{S}^{1}\text{ and }\frac{1}{2}\mathbbm{S}^{1}\right]$
$\leq\mathbbm{P}[\mathrm{CR}(\rho,0)<2\sqrt{\delta}]\cdot S+(4\sqrt{\delta})^{\frac{1}{4}+o(1)}.$ (4.12)
Finally, by combining (4.8), (4.9), (4.10), (4.12), we have (recall that $\delta=2\sqrt{\varepsilon}$)
$\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]\leq\mathbbm{P}[\mathrm{CR}(\rho,0)<2\sqrt{\delta}]\cdot S+(4\sqrt{\delta})^{\frac{1}{4}+o(1)}+C_{1}\varepsilon^{\frac{1}{4}}+2C_{1}S\varepsilon^{\frac{1}{8}}.$
Using Corollary 3.7, we conclude that $\limsup_{\varepsilon\to 0}\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]\cdot(\log|\log\varepsilon|)\leq 3S$, as desired. ∎